Vector a has an x component of 60 n and a y component of 0 n. Vector b has an x component of -45 n and a y component of 75 n. What is the x component of a-b? What is the y component of a-b?
1
Expert's answer
2018-01-24T09:38:08-0500
Given: a (60, 0); b (-45, 75)
x-component = 60 - (-45) = 105
y-component = 0 - 75 = -75